\begin{thebibliography}{99}

\bibitem{BD}
G. Bell and A. N. Dranishnikov, 
{\it On asymptotic dimension of groups},
Algebraic and Geometric Topology 1 (2001),  57--71.

\bibitem{BD1} 
G. Bell and A. N. Dranishnikov, 
{\it Asymptotic dimension}, preprint on the Archives http://arxiv.org/abs/math/0703766v2

\bibitem{D2}  
A. N. Dranishnikov, 
{\it On the theory of extensions of mappings of compact spaces,} (Russian)
Uspekhi Mat. Nauk {\bf 53}:5 (1998), 65--72;
English transl. in Russian Math. Surveys {\bf 53}:5 (1998), 929--935.

\bibitem{D} A. N. Dranishnikov, 
{\it Asymptotic topology,} Russian Math. Surveys {\bf 55}(2000), No 6, 71--116.

\bibitem{DD2} 
A. Dranishnikov and J. Dydak,
{\it Extension dimension and extension types},
Trudy Mat. Inst. Steklova {\bf 212} (1996), Otobrazh. i Razmer., 61--94; 
English transl. in Proc. Steklov Inst. Math. {\bf 212}:1 (1996), 55--88.

\bibitem{DD} 
A. Dranishnikov and J. Dydak,
{\it Extension theory of separable metrizable spaces with applications to dimension theory},
Trans. Amer. Math. Soc. {\bf 353}:1 (2001) 133--156.

\bibitem{G}
M. Gromov,  
{\it Asymptotic invariants of infinite groups},
Geometric group theory, Vol. 2 (Sussex, 1991), London Math. Soc. Lecture Note Ser., vol. 182, Cambridge Univ. Press, Cambridge, 1993, pp. 1–-295.

\bibitem{M} 
P.  Mitchener, 
{\it Coarse homology theories},
Algebr. Geom. Topol. {\bf 1}  (2001), 271--297 (electronic).

\bibitem{R1} 
J. Roe, 
{\it From foliations to coarse geometry and back},
in: Analysis and geometry in foliated manifolds, (X. Masa, E. Macias-Virg\'os, J.A.Alvarez L\'opez, Editors), World Scientific, Singapore 1995, pp. 195-–206.

\bibitem{R}
J. Roe,
{\it Warped cones and property A},
Geometry $\&$ Topology {\bf 9} (2005) 163–178 
Published: 6 January 2005, Corrected: 7 March 2005 

\bibitem{S} 
M. Sawicki,
{\it Absolute extensors and absolute neighborhood extensors in asymptotic categories},
Topology Appl. {\bf 150}:1–3 (2005), 59-–78.

\end{thebibliography}

\documentclass[amsmath,amscd,amsbsy,amssymb]{amsart}
 \relax

\chardef\bslash=`\\ % p.  424, TeXbook %\newcommand{\ntt}{\seriesm\shape n\tt}

\makeatletter \def\verbatim{\interlinepenalty\@M \@verbatim
\leftskip\@totalleftmargin\advance\leftskip2pc \frenchspacing\@vobeyspaces
\@xverbatim} \makeatother \hfuzz1pc
\makeatletter \def\dgt@k{\dg@DX=-3 \dg@DY=2 \dg@SIZE=3} \makeatother
\makeatletter \def\dgt@kk{\dg@DX=3 \dg@DY=-1 \dg@SIZE=3}% \makeatother
\theoremstyle{plain} \newtheorem{thm}{Theorem}[section]
\newtheorem{cor}[thm]{Corollary} \newtheorem{lemma}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\theoremstyle{definition} 
 \newtheorem{que}[thm]{Question}
\newcommand{\A}{\mathcal A}
\newcommand{\LL}{\mathcal L}

\numberwithin{equation}{section}
\newcommand{\aed}{{\mathrm a}{\mathrm s} {\mathrm -}{\mathrm e}{\mathrm x}
{\mathrm t} {\mathrm -}{\mathrm d}{\mathrm i}{\mathrm m}}
\newcommand{\ed}{{\mathrm e}{\mathrm x}
{\mathrm t}{\text -}{\mathrm d}{\mathrm i}{\mathrm m}}
\newcommand{\asd}{{\mathrm a}{\mathrm s} {\mathrm d}{\mathrm i}{\mathrm m}}
\newcommand{\Var}{{\mathrm V}{\mathrm a}{\mathrm r}}
\newcommand{\R}{{\mathbb R}}
\newcommand{\OO}{{\mathcal O}}
\newcommand{\N}{{\mathbb N}}

\newcommand{\diam}{{\mathrm d}{\mathrm i} {\mathrm a}{\mathrm m}\ }

\usepackage[all]{xy} %\usepackage[ukrainian]{cyrtext}
\begin{document}

%%%%%%% Begin Topmatter %%%%%%%%%%
\title[On asymptotic extension dimension]{On asymptotic extension dimension}

\author{Du\v{s}an Repov\v{s}}
\address{Faculty of Mathematics and Physics, and Faculty of Education, 
University of Ljubljana,
Jadranska 19, 1000 Ljubljana, Slovenia} \email{dusan.repovs@guest.arnes.si}

\author{Mykhailo Zarichnyi}
\address{Department of Mechanics and Mathematics, Lviv National University,
Universytetska 1, 79000 Lviv, Ukraine} \email{mzar@litech.lviv.ua}

\subjclass[2010] {Primary: 54F45, 54C40; Secondary: 55M10}

\keywords{Asymptotic category, 
asymptotic extension dimension, 
Higson compactification, 
Higson corona, 
coarse uniform map, 
coarse CW complex, 
proper asimptotically Lipschitz map,
slowly oscillating map,
almost geodecis space,
Lipschitz neighborhood euclidean extensor, 
foliated warped cone}

\begin{abstract}{The aim of this paper  is to introduce an asymptotic
counterpart of the extension dimension defined by Dranishnikov. The
main result establishes a relation between the asymptotic
extensional dimension of a proper metric space and extension
dimension of its Higson corona.}
\end{abstract}

\date{\today}

\thanks{This research was supported by the
Slovenian Research Agency grants P1-0292-0101 and J1-2057-0101.}

\maketitle

\section{Introduction}\label{S:intro}

Asymptotic dimension defined by Gromov \cite{G} has been
an object of
study in numerous publications (see expository paper
\cite{BD1}).
A metric space $(X,d)$ is of asymptotic dimension $\le n$ (written
$\mathrm{asdim}\,X\le n$) if for every $D>0$ there exists a
uniformly bounded cover $\mathcal U$ of $X$ such that $\mathcal
U=\mathcal U^0\cup\dots\cup\mathcal U^n$, where every family
$\mathcal U^i$ is $D$-disjoint, $i=0,1,\dots,n$. Recall that a
family $\mathcal A$ of subsets of $X$ is {\em uniformly bounded}
if $$\mathrm{mesh}\,\mathcal A=\sup\{\mathrm{diam}\,A\mid
A\in\mathcal A\}<\infty$$ and is called $D$-{\em disjoint} if
$\inf\{d(a,a')\mid a\in A,\ a'\in A'\}>D$, for every distinct
$A,A'\in \mathcal A$.

The asymptotic dimension can be characterized in different terms;
in particular, in terms of extension of maps into euclidean spaces.
The aim of this paper is to introduce an asymptotic analogue of
the extension dimension introduced by Dranishnikov in \cite{D2,
D}.

\section{Preliminaries}\label{S:prelim}

A typical metric is denoted by $d$. By $N_r(x)$ we denote the open ball of
radius $r$ centered at a point $x$ of a metric space.

\subsection{Asymptotic category}
 A map $f\colon X\to Y$ between metric
spaces is called $(\lambda,\varepsilon)$-{\em Lipschitz\/} for $\lambda>0$,
$\varepsilon\ge0$ if $d(f(x),f(x'))\le\lambda d(x,x')+\varepsilon$ for every
$x,x'\in X$. A map is called {\em asymptotically Lipschitz\/} if it is
$(\lambda,\varepsilon)$-Lipschitz for some $\lambda,\varepsilon>0$.
The $(\lambda,0)$-Lipschitz maps are also called {\em $\lambda$-Lipschitz\/},
$(1,0)$-Lipschitz maps are also called {\em short}.
A metric space $X$ is called {\em proper\/} if every closed ball in $X$
is compact.

The {\em asymptotic category\/} $\A$ was introduced by Dranishnikov \cite{D}.
The objects of $\A$ are proper metric spaces and the morphisms are proper
asymptotically Lipschitz maps. Recall that a map is called {\em proper} if the
preimage of every compact set is compact.

We also need the notion of a {\em coarse map}. A map between proper metric
spaces is called {\em coarse uniform\/} if
for
every $C>0$ there is $K>0$ such that for every $x,x'\in X$ with $d(x,x')<C$ we
have $d(f(x),f(x'))<K$. A map $f\colon X\to Y$  is called {\em metric proper\/}
if the preimage $f^{-1}(B)$ is bounded for every bounded set $B\subset Y$.
A map is {\em coarse} if it is metric proper and coarse uniform.

\subsection{Higson compactification and Higson corona}

Let $\varphi\colon X\to\R$ be a function defined on a metric space $X$.
For every $x\in X$ and every $r>0$ let
$V_r\varphi(x)=\sup\{|\varphi(y)-\varphi(x)|\mid
y\in N_r(x)\}$. A function $\varphi$ is called {\it slowly oscillating\/}
whenever for every $r>0$ we have
$V_r\varphi (x)\to0$ as $x\to\infty$ (the latter means that for every
$\varepsilon>0$
there exists a compact subspace $K\subset X$ such that
$|V_r\varphi(x)|<\varepsilon$
for all $x\in X\setminus K$. 

Let $\bar X$ be the compactification of $X$ that
corresponds
to the family of all continuous bounded slowly oscillation functions. The {\it
Higson
corona\/} of $X$ is the remainder $\nu X=\bar X\setminus X$ of this
compactification.
It is well-known that the Higson corona is a functor from the category of proper
metric space and coarse maps into the category of compact Hausdorff spaces. In
particular, if $X\subset Y$, then $\nu X\subset \nu Y$.

For any subset $A$ of $X$ we denote by $A'$ its trace on $\nu X$, i. e. the
intersection of the closure of $A$ in $\bar X$ with $\nu X$. Obviously, the set
$A'$ coincides with the Higson corona $\nu A$.

\subsection{Cones}

Let $X$ be a metric space of diameter $\le1$. The {\em open cone\/}
of $X$ is the set $\OO X=(X\times\R_+)/(X\times\{0\})$ endowed with
the metric (by $[x,t]$ we denote the equivalence class of $(x,t)\in
X\times\R_+$):
$$d([x_1,t_1],[x_2,t_2])=|t_1-t_2|+\min\{t_1,t_2\}d(x_1,x_2).$$
For a map $f\colon X\to Y$ of metric spaces we denote by $\OO f\colon \OO X\to
\OO Y$ the
map
defined as $\OO f([x,t])=[f(x),t]$.

\begin{prop} If $f\colon X\to Y$ is a Lipschitz map than $\OO f$ is an
asymptotically
Lipschitz map.
\end{prop}
\begin{proof} Suppose a map $f\colon X\to Y$ is $\lambda$-Lipschitz. Then for
any $[x_1,t_1], [x_2,t_2]\in \OO X$ we have
\begin{align*}d(\OO f([x_1,t_1]), \OO f([x_2,t_2]))=&d([f(x_1),t_1],
[f(x_2),t_2])\\=&|t_1-t_2|+\min\{t_1,t_2\}d(f(x_1),f(x_2))\\ \le& \lambda'
(|t_1-t_2|+\min\{t_1,t_2\}d(x_1,x_2)),
\end{align*}
where $\lambda'=\max\{\lambda,1\}$.
\end{proof}
The open cone of a finite CW-complex is a coarse CW-complex in the sense of
\cite{M}.

 Denote by $\alpha_L\colon \OO L\to\R$ the function defined by
$\alpha_L([x,t])=t$.
Obviously, $\alpha_L$ is a short function.
Let $\tilde \OO L=\{[x,t]\in \OO L\mid t\ge1\}$. Denote by $\beta_L\colon \tilde
\OO L\to
L$ the map $\beta_L([x,t])=x$.
\begin{lemma} The map $\beta_L$ is slowly oscillating.
\end{lemma}
\begin{proof} For $R>0$, the $R$-ball centered at $[x,0]$ is $\{[x,t]\mid t<R$.
If $d([x,t],[x_1,t_1])<K<R$, then $|t-t_1|+\min\{t,t_1\}d(x,x_1)<K$, i.e.
$(t-R)d(x,x_1)<R$ and $d(x,x_1)<K/(t-K)$. Therefore,
$d(\beta_L(x),\beta_L(x_1))<K/(R-K)\to 0$ as $R\to\infty$.
\end{proof}

Let $\bar\beta_L\colon \tilde \OO L\to L$ be the (unique) extension of the map
$\beta_L$. Denote by $\eta_L\colon \nu\tilde \OO L\to L$ the restriction of
$\beta_L$.
\begin{prop} Let $f\colon A \to \OO L$ be a proper asymptotically Lipschitz map
defined on a proper closed subset $A$ of a proper metric space $X$. There exists
a
neighborgood $W$ of $A$ in $X$, a  proper asymptotically Lipschitz map $g\colon
W\to \OO L$ with the following property: there exist constants $\lambda,s>0$
such
that $\alpha_L(g(a))\le \lambda d(a, X\setminus W)+s$.
\end{prop}
\begin{proof} We may assume that $L$ is a subset of $I^n$ for some $n$ and there
exists a Lipschitz retraction $r\colon U\to L$ of a neighborhood $U$ of $L$ in
$I^n$. Since $\OO I^n$ is Lipschitz equivalent to $\R^{n+1}_+$, there exists a
$(\lambda',s')$-Lipschitz extension $\tilde g\colon X\to \OO I^n$ of $g$.

Put $W=\tilde g^{-1}(\OO U)$ and $\bar g=\tilde g|W$. For every $a\in A$ and
$w\in
X\setminus W$ we have $$d(g(a),\tilde g(w)\le \lambda' d(a,w)+s'\le
\lambda'd(a,X\setminus W)+s.$$

Suppose that $d(L,I^n\setminus U)=c>0$, then, since $\tilde g(w)\notin CU$,
\begin{align*} d(g(a),\tilde g(w)= &|\alpha_L(g(a))-\alpha_L(\tilde g(w))| \\
+&\min\{\alpha_L(g(a)),\alpha_L(\tilde g(w))\} d(\beta_L(g(a)), \beta_L(\tilde
g(w))) \\
\ge & |\alpha_L(g(a))-\alpha_L(\tilde
g(w))|+c\min\{\alpha_L(g(a)),\alpha_L(\tilde g(w))\}\\
\ge & c'\alpha_L(g(a)),
\end{align*}
where $c'=\min\{c,1\}$. Then $\alpha_L(g(a))\le \lambda d(a, X\setminus W)+s$,
where $\lambda =\lambda'/c'$, $s=s'/c'$.\end{proof}

\section{Auxiliary results}

In this section we shall
collect some results needed for the proof of the
main result. They are proved in \cite{D} but it turns out that we
have also to cover the case of functions with infinite values.

A map $f\colon X\to\R_+\cup\{\infty\}$ is said to be {\em coarsely proper\/} if
the preimage $f^{-1}([0,c])$ is bounded for every $c\in \R_+$.
\begin{lemma}\label{l:1} For any function $\varphi\colon X\to  \R_+$ with
$\varphi(x)\to0$
as $x\to\infty$ the function $1/\varphi\colon X\to  \R_+\cup\{\infty\}$ is
coarsely proper.
\end{lemma}
\begin{prop}\label{p:1} Let $f\colon X\to\R_+\cup\{\infty\}$ be a coarsely
proper function. There exists an asymptotically Lipschitz proper function
$q\colon X\to\R_+$ with $q\le f$.
\end{prop}
\begin{proof} This was proved in \cite{D} for the case of $f\colon X\to\R_+$
(see Proposition 3.5). That proof also works in our case.
\end{proof}
\begin{prop}\label{p:2} Let $f_n\colon X\to\R_+\cup\{\infty\}$ be a sequence of
coarsely
proper functions. Then there exists a filtration $X=\cup_{n=1}^\infty A_n$ and a
coarsely proper function $f\colon X\to \R_+$ with $f|A_n\le n$ and
$f|(X\setminus A_n)\le f_n$ for every $n$.
\end{prop}
\begin{proof} Let $B_n=\cup_{i=1}^nf^{-1}_i([0,n])$. the sets $B_i$ are bounded
and $B_1\subset B_2\subset\dots$. Therefore, there exist bounded subsets
$A_1\subset A_2\subset\dots$ such that $A_n\cap(\cup_{i=1}^\infty B_i)=B_n$ and
$\cup_{i=1}^\infty A_i=X$. For $x\in A_n\setminus A_{n-1}$, put $f(x)=n$.
Obviously, $f$ is coarsely proper and $f|A_n\le n$. Now suppose that $x\notin
A_n$, then  $x\notin B_n$ and therefore $x\notin f^{-1}_n([0,n])$, i.e.
$f_n(x)>n\ge f|(X\setminus A_n)$.
\end{proof}

The following is an easy modification of Lemma 3.6 from \cite{D} and the proof
of it works in our case as well.
\begin{lemma} Suppose that $f\colon A\to \R_+\cup\{\infty\}$ is a coarsely
proper map defined on a closed subset $A$ of a proper metric space $X$ and
$g\colon W\to \R_+$ is a proper asymptotically Lipschitz map such that $g\le
f|W$ and  there exist $\lambda,s$ such that $\lambda d(a,X\setminus W)+s\ge
g(a)$ for every $a\in A$. Then there exists a proper  asymptotically Lipschitz
map $\bar g\colon X\to \R_+$ for which $\bar g\le f$ and $\bar g|A=g$.
\end{lemma}

\subsection{Almost geodesic spaces}
A metric space $X$ is said to be {\em almost geodesic\/} if there exists $C>0$
such that  for every two points $x,y\in X$ there is a short map  $f\colon
[0,Cd(x,y)]\to X$
with $f(0)=x$, $f(Cd(x,y))=y$. If in this definition $C=1$, then we come to the
well-known notion of {\em geodesic} space.

We are going to describe a construction of embedding of a discrete metric space
$X$ into an almost geodesic space of the  asymptotic dimension $\min\{\asd
X,1\}$.
For an unbounded  discrete metric space $X$ with base point
$x_0$  define a function $f\colon X\to[0,\infty)$ by the formula
$f(x)=d(x,x_0)$. Choose a sequence $0=t_0<t_1<t_2<\dots$ in $f(X)$ so that
$t_{i+1}>2t_i$ for every $i$. To every pair of points $x,y\in
f^{-1}([t_i,t_{i+1}])$, for some $i$, attach the line segment $[0,d(x,y)]$ along
its endpoints. Let $\hat X$ is the union of $X$ and all attached segments. We
endow $\hat X$ with the
maximal metric that agrees with the initial metric on $X$ and the standard
metric on every attached segment.

Note that since $X$ is discrete and proper, every set $f^{-1}([t_i,t_{i+1}])$ is
finite and therefore  $\hat X$ is a proper metric space.

\begin{prop} The space $\hat X$ is almost geodesic.
\end{prop}
\begin{proof}
Suppose that $x,y\in
\hat X$, then
$x\in [x_1,x_2]$, $y\in [y_1,y_2]$, where $x_1,x_2,y_1,y_2\in X$ and
$[x_1,x_2]$, $[y_1,y_2]$ are attached segments. We may suppose that
$d(x,y)=d(x,x_1)+d(x_1,y_1)+d(y_1,y)$.

Case 1): There exists $i$ such that $x_1,y_1\in f^{-1}([t_i,t_{i+1}])$. Then
$[x,x_1]\cup[x_1,y_1]\cup[y_1,y]$ is a segment of diameter $d(x,y)$ that
connects $x$ and $y$ in $\hat X$.

Case 2): $f(x_1)\in[t_i,t_{i+1}]$, $f(y_1)\in[t_j,t_{j+1}]$, where $i\neq j$.
Without loss of generality, we may assume that $i<j$.

Obviously, $d(x_1,y_1)\le d(x,y)$. Since $|t_{j}-t_{j-1}|\le d(x_1,y_1)$, we see
that  $|t_{j}-t_{j-1}|\le d(x,y)$. This implies that $t_j/2\le d(x,y)$, or
equivalently, $t_j\le d(x,y)$.
Besides, $d(y_1, f^{-1}([0,t_{j-1}])))\le d(x_1,y_1)\le d(a,b)$.

For every $k=i,i+1,\dots,j_1$ choose $z_k\in f^{-1}(t_k)$. Then
\begin{align*} d(y_1,z_{j-1})\le &d(y_1,f^{-1}([0,t_{j-1}]) +\diam
(f^{-1}([0,t_{j-1}]))\\ \le &
d(a,b)+2t_{j-1}\le d(a,b)+t_{j}\le 3d(a,b).
\end{align*}
We connect $x$ and $y$ by the segment
$$J=[x,x_1]\cup[x_1,z_1]\cup\cup_{k=i}^{j-1}[z_k,z_{k+1}])
\cup[z_{j-1},y_1]\cup[y_1,y].$$
Then
\begin{align*} \diam J\le &
d(x,x_1)+d(x_1,z_{i+1})+\left(\sum_{k=i+1}^{j-1}d(z_k,z_{k+1})\right)+
d(z_{j-1},y_1)+d(y_1,y)\\=& d(x,y)+2t_{i+1}+
\sum_{k=i+1}^{j-1}2t_{k+1}+5d(x,y)+d(x,y)\\ \le& 7d(x,y)+2(t_{i+1}+\dots+t_j)\le
7d(x,y)+4t_j\\ \le&15d(x,y).
\end{align*}
\end{proof}
We need a version of the fact proved in \cite{D} for geodesic spaces.

\begin{prop}\label{p:geod} Let $f\colon X\to Y$ be a coarse uniform map of an
almost geodesic
space
$X$. Then $f$ is asymptotically Lipschits.
\end{prop}
\begin{proof} Let $C$ be a constant from the definition of almost geodesic
space. Suppose $x,y\in X$, then there exists a short map $\alpha\colon [0,
Cd(x,y)]\to X$ such that $\alpha(0)=x$, $\alpha(Cd(x,y))=y$. There exist points
$0=t_0<t_1<\dots<t_{k-1}<t_k=Cd(x,y)$, where $k\le[d(x,y)]+1$, such that
$|t_i-t_{i-1}|\le C$ for every $i=1,\dots,k$.

Since $f$ is coarse uniform, there exists $R>0$ such that $d(f(x'),f(y')<R$
whenever $d(x',y')\le C$. Then
\begin{align*}d(f(x),f(y))\le&\sum_{i=1}^kd(f(\alpha(t_i)),
f(\alpha(t_{i-1})))\le kR\le ([d(x,y)]+1)R\\ \le& Rd(x,y)+2R.
\end{align*}
\end{proof}

\section{Asymptotic extension dimension}

Let $P$ be an object of the category $\A$. For any object $X$ of
$\A$ the {\em Kuratowski notation\/} $X\tau P$ means the following:
for every proper asymptotically Lipschitz map $f\colon A\to P$
defined on a closed subset $A$ of $X$ there is a proper
asymptotically Lipschitz extension of $f$ onto $X$.

Denote by $\LL$ the class of compact absolute Lipschitz neighborhood euclidean
extensors (ALNER).
Following \cite{}, we define a preorder relation $\leq$ on $\LL$. For
$L_1,L_2\in\LL$, we have $L_1\leq L_2$ if and only if $X\tau \OO L_1$ implies
$X\tau \OO L_2$ for all proper metric spaces $X$. This preorder relation leads
to the following equivalence relation $\sim $ on $\LL$: $L_1\sim L_2$ if and
only if $L_1\leq L_2$ and $L_2\leq L_1$. We denote by $[L]$ the equivalence
class containing $L\in\LL$.

For a proper metric space $X$, we say that its {\em asymptotic
extension dimension does not exceed $[\OO L]$} (briefly $\aed X\le
[\OO L]$ whenever $X\tau\OO L$.
If $\aed X\le [\OO L]$, then the equality $\aed X= [\OO L]$ means
the following. If we also have $\aed X\le [\OO L']$, then  $[\OO
L] \le[\OO L']$.
By  \cite{D} (see also \cite{S}), the element $[*]$
is maximal.

\begin{thm} Let $L$ be a compact metric ALNER. The following conditions are
equivalent:
\begin{enumerate}
\item $\aed X\le[\OO L]$;
\item $\ed \nu X\le[L]$.
\end{enumerate}
\end{thm}
\begin{proof} Assume that $\aed X\le[\OO L]$. Let $\varphi\colon C\to L$ be a
map
defined on a closed subset $C$ of $\nu X$. Since $L\in \mathrm{ANE}$, there
exists an
extension $\varphi'\colon V\to L$ of $\varphi$ over a closed neighborhood $V$ of
$C$ in $\bar X=X\cup\nu X$. Then $\Var_R\varphi'(x)\to0$ as $x\to\infty$, for
any fixed $R>0$. By Lemma \ref{l:1}, the function $$f_n\colon V\cap
X\to\R_+\cup\{\infty\},\
f_n(x)=\frac{1}{\Var_R\varphi'(x)},$$
is coarsely proper, for every $n\in\N$. By Proposition \ref{p:2}, there is a
coarsely proper function $f\colon V\cap X\to\R_+$ and a filtration $V\cap
X=\cup_{n=1}^\infty A_n$ such that $f|A_n\le n$ and $f|(X\setminus A_n)\le f_n$.
By Proposition 3.5 from \cite{D}, there is an asymptotically Lipschitz function
$q\colon
V\cap X\to\R_+$ with $q\le f$. We suppose that $q$ is $(\lambda,s)$-Lipschitz
for some $\lambda,s>0$. Define the map $g\colon V\cap X\to \OO L$ by the
formula $g(x)=[\varphi'(x),q(x)]$.

We are going to check that the map $g(x)$ is asymptotically Lipschitz. Let
$x,y\in V\cap X$ and $n-1\le d(x,y)\le n$.

Suppose that $x,y\in(V\cap X)\setminus A_n$, then $q(x)\le f_n(x)$, $q(y)\le
f_n(y)$. We have
\begin{align*}
d(g(x),g(y))=&|q(x)-q(y)|+\min\{q(x),q(y)\}d(\varphi'(x),\varphi'(y))\\
\le& \lambda d(x,y)+s+\min\{q(x),q(y)\}\Var_n\varphi'(x)\\
\le& \lambda d(x,y)+s+1.
\end{align*}

If $x\in A_n$, then $q(x)\le n$ and we obtain
\begin{align*} d(g(x),g(y))\le& \lambda d(x,y)+s+nd(\varphi'(x),\varphi'(y))\\
\le& \lambda d(x,y)+s+n\diam L\le \lambda d(x,y)+s+(d(x,y)+1)\diam L\\
\le&(\lambda+\diam L)d(x,y)+(s+\diam L).
\end{align*}
We argue similarly if $y\in A_n$.
Now, by the assumption, there is an asymptotically Lipschitz extension $\bar
g\colon X\to \OO L$ of $g$. Consider the composition $\eta_L\nu \bar g\colon \nu
X\to \OO L$. Obviously,  $\eta_L\nu \bar g|C=\varphi$.

Let $f\colon A\to \OO L$ be an asymptotically Lipschitz map defined on a proper
closed subset $A$ of a proper metric space $X$. By Proposition 2.2,
there is
a proper  asymptotically Lipschitz map $\tilde f\colon W\to \OO L$ and constants
$\lambda, s$ such that $\alpha_Lf(a)\le\lambda d(a,X\setminus W)+s$ for all
$a\in A$. Denote by
$\varphi\colon \nu X\to L$ an extension of the composition $\eta_L\nu \tilde f$.
Since
$L$ is an absolute neighborhood extensor, there exists an extension $\psi\colon
V\to L$ of $\varphi$ onto a closed neighborhood of $\nu X$ in the Higson
compactification $\bar X$. Extend $\psi$ to a map  $\hat\psi\colon(V\cap
X)\hat{}\to L$ as follows. Let
$J$ be a segment attached to $V$ with endpoints $a$ and $b$. We require that
$\hat\psi$ linearly maps $J$ onto a geodesic segment in $L$ with endpoints
$\psi(a)$
and $\psi(b)$.

We show that $\hat\psi $ is a slowly oscillating map. Since $\psi$ is slowly
oscillating, for every $\varepsilon>0$ and  $R>0$ there exists $K>0$ such that
$\Var_R\psi(x)<\varepsilon$ whenever $d(x,x_0)>K$. Suppose that $\hat\psi$ is
not slowly oscillating, then there exist $R>0$, $C>0$,  and sequences $(x_1^i)$,
$x_2^i$  in $(V\cap X)\hat{}$ such that $d(x_1^i,x_2^i)<R$,  $x_1^i\to\infty$,
$x_2^i\to\infty$ and $d(\hat\psi(x_1^i),\hat\psi(x_2^i))>C$ for every $i$. We
assume that $x_1^i\in[a_1^i,b_1^i]$, $x_2^i\in[a_2^i,b_2^i]$, for every $i$,
where $a_1^i,b_1^i,a_2^i,b_2^i\in X\cap V$.  Without loss of generality we may
assume that $a_1^i\to\infty$ and there exists $C_1>0$ such that
$d(\hat\psi(x_1^i),\hat\psi(a_1^i))>C_1$ for every $i$. If $d(a_1^i,b_1^i)<K$
for all $i$ and some $K>0$, then
$d(\hat\psi(x_1^i),\hat\psi(a_1^i))<d(\hat\psi(a_1^i),\hat\psi(b_1^i))\to0$, and
we obtain a contradiction. Therefore, we may assume that  $d(a_1^i,b_1^i)\to
\infty$. Then $d(a_1^i,x_1^i)/d(a_1^i,b_1^i)<R/d(a_1^i,b_1^i)\to0$ and
therefore, by the definition of the map $\hat\psi$,
$d(\hat\psi(x_1^i),\hat\psi(a_1^i))/d(\hat\psi(a_1^i),\hat\psi(b_1^i))\to0$.
Then obviously $d(\hat\psi(x_1^i),\hat\psi(a_1^i))\to0$ and we obtain a
contradiction.

Since the map $\tilde f$ is asymptotically Lipschitz, there exists
$K>0$ such that for any $a\in W$ we have $$\diam ( \alpha_L\tilde
f(N_1(a))+  \alpha_L\tilde f(a)\diam(\psi(N_1(a))\le K.$$ Define
the function $r\colon (X\cap V)\hat{}\to\R_+\cup\{\infty\}$ by the
formula $r(x)= K/(\psi(N_1(x)))$. We have $ f(a)\le r(a)$ for
every $a\in A$. The function $r$ is asymptotically proper and by
Proposition \ref{p:1}, there exists a $(\lambda',s')$-Lipschitz
function $\bar f\colon X\to\R_+$, for some $\lambda',s'$,  with
$\bar f\le r$ and $\bar f|A=\alpha_Lf$.

Define a map $g\colon (X\cap V)\hat{}\to\R$ by the
formula $g(x)=(\psi(x),\bar f(x))$. Obviously, $g|A=f$. We are going to show
that $g$ is a coarse uniform map.
Suppose $x,y\in X$, $d(x,y)<1$, then
\begin{align*}d(g(x), g(y))\le &|\bar f(x)-\bar f(y)|+\min\{\bar
f(x),\bar f(y)\}d(\psi(x),\psi(y)) \\ \le\lambda'+s'+K.\end{align*}

Note that, since $\bar f$ is proper, $g$ is also proper. Since $g$ is coarse
uniform, by Proposition \ref{p:geod}, $g$ is asymptotically Lipschitz.
\end{proof}
\begin{cor} {\bf (Finite Sum Theorem)} Suppose $X$ is a proper metric space,
$X=X_1\cup X_2$, where $X_1, X_2$ are closed subsets of $X$ with
$\aed X_i\le[\OO L]$, $i=1,2$, for some $L\in\LL$. Then $\aed
X\le[\OO L]$.
\end{cor}
\begin{proof} Since $\nu X=\nu X_1\cup \nu X_2$, the result follows from Theorem 4.1
and the finite sum theorem for extension dimension (see \cite{BD1}).
\end{proof}

\section{Remarks and open questions}

\begin{que} Does the equality $\aed \mathbb R^n=S^n$ hold?
\end{que}

\begin{que} Let $L_1,L_2$ be finite polyhedra in euclidean spaces endowed with
the induced metric. Is the inequality $[L_1]\le[L_2]$ introduced
in \cite{D2} equivalent to the inequality $[L_1]\le[L_2]$ as in
Section 4? 
\end{que}

One can define analogue of the asymptotic extension dimension
by using warped cones instead of open cones. Following \cite{R1}
we review this construction briefly. Let $\mathcal F$ be a
foliation on a compact smooth manifold $V$. Let $N$ be any
complementary subbundle to $T\mathcal F$ in $TM$.  Choose
Euclidean metrics $g_N$ in $N$ and $g_{\mathcal F}$ in $T\mathcal
F$. The {\em foliated warped cone} ${\mathcal O}_{\mathcal F}$ is
the manifold $V \times [0,\infty)/V\times\{0\}$ equipped with the
metric induced for $t \ge 1$ by the Riemannian metric $g_R +
g_{\mathcal F} + t2g_N$. The metric structure on any bounded
neighborhood of the distinguished point is irrelevant.

\begin{que} Is the obtained warped cone an absolute neighborhood extensor in the asymptotic category?
\end{que}

An affirmative answer to this question would allow us to define
asymptotic extension dimension theory with the values in warped
cones.

%\section{Acknowledgements}

\end{document}